• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Zhihong Ye

    Log entry time 09:38:28 on March 09, 2012

    Entry number 365066

    keyword=G2p Status Update

    Everything was running very smooth yesterday. 
    We finished the Harp scan, although it is still noisy, and calibrated BPM
    and slow raster.
    All optics data at this configuration have been taken and part of data
    for dilution study were taken before we open up the hall this morning. 
    Now we are moving the Chicane to its new location (Conf#8). And after it
    finished, we will ramp up target magnet to 5T. The hall will be close
    about at 8pm tonight.
    Experts continues working on diagonize the harp problem. One thing they
    found out is that the filter they put it (before they found out the
    grouding issue) was not doing too much help but instead introduct another
    kind of wave-form noise. They took that out and will see whether it will
    help or not when beam come back.
    After beam permit and MCC finishes beam tuning, we will do harp scan, 
    bpm & raster calibration again then take optics data during owl shift if
    we have time. Tomorrow during the day shfit, we are going to polarize the
    target and start taking Gep production data.

    A copy of this log entry has been emailed to: jpchen,slifer,camsonne,doug,rom,brads